OpenStudy (anonymous):

Find the vertex, focus, directrix, and focal width of the parabola. x = 4y^2

OpenStudy (perl):

The general form of your parabola is $$ \Large x = \frac {1}{4p} y^2 $$Focal width is 4p

OpenStudy (anonymous):

ok what do i do wuth that formula

OpenStudy (perl):

we need to solve 1/(4p) = 4

OpenStudy (anonymous):

oh so p = 1/16?

OpenStudy (perl):

yes

OpenStudy (perl):

the vertex is (h,k) directrix is x = -p focal width is 4p

OpenStudy (perl):

it is centered at the origin

OpenStudy (anonymous):

how do we find h and k

OpenStudy (anonymous):

ok the focal width is .25

OpenStudy (perl):

The general form of a parabola that opens to the right or left \[ \Large x-h = \frac {1}{4p} (y-k)^2 \]Vertex is (h,k) directrix is x =h -p focal width is 4p

OpenStudy (perl):

yes that is correct

OpenStudy (anonymous):

is the focal width represented as a single letter?

OpenStudy (perl):

4 times that value of p

OpenStudy (anonymous):

like i know i can plug in x as 4y^2 but thats still not enough to find either h or k

OpenStudy (perl):

h must be zero and k must be zero

OpenStudy (anonymous):

how do you know its at the origin though

OpenStudy (anonymous):

oh wait all the choices have the vertix set to (0,0) but if it isnt how do you know

OpenStudy (perl):

\[ \Large x-h = \frac {1}{4p} (y-k)^2 \iff x = \frac {1}{4p} y^2 \]

OpenStudy (perl):

the two equation left sides must match and the right side must match

OpenStudy (anonymous):

ohhhhhh omg ok i get it

OpenStudy (perl):

the two equations left sides must match and right sides

OpenStudy (anonymous):

and thats how you know its at the origin?

OpenStudy (perl):

yes. or by graphing it

OpenStudy (anonymous):

ok then the directrix is 0-.25 which would = -1/4

OpenStudy (anonymous):

but what about the focus?

OpenStudy (perl):

correct. and the focus is (h+ p, k )

OpenStudy (anonymous):

oh ok this is actually simplier than i thought! thank you so much!